# An Online Learning Approach to Prompt-based Selection of Generative Models

### v2

@@ 6.2 Summary
The main finding of our numerical experiments is the improvement of the proposed contextual bandit PAK-UCB algorithm over the one-arm oracle baseline. This result means that the online learning algorithm can outperform a user with side-knowledge of the single best-performing model, which is made possible by a prompt-based selection of the model. This finding supports the application of contextual bandit algorithms in the selection of text-based generative models. Moreover, our numerical results indicate that the proposed PAK-UCB algorithm can significantly outperform the standard LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) methods, showcasing the effectiveness of our framework in the prompt-based model selection setting. Additionally, the proposed PAK-UCB algorithm can perform better with a non-linear kernel function. Finally, in our experiments, the proposed RFF-UCB variant could reduce the computational costs of the general PAK-UCB algorithm.
6.3 Experimental Setups
@@
Setup 1. Prompt-based selection among standard text-to-image models. The first set of experiments are on the setup illustrated in Figure 1(a), where we generate images from Stable Diffusion v1.5 and PixArt-$\alpha$. The results show that PAK-UCB-poly3 outperforms the baseline algorithm and attains a high optimal-pick-ratio, which shows that it can identify the optimal model conditioned to the prompt (see Figure LABEL:fig:t2itab1). Additionally, we provide numerical results on various T2I generative models, including UniDiffuser and DeepFloyd (see Figure LABEL:fig:t2itab2).
@@
Setup 2. Adaptation to new prompt types and models. We consider scenarios where new text-to-image models or prompt types are introduced after the initial deployment. At the beginning of the first experiment, Stable Diffusion v1.5 and PixArt-$\alpha$ are available, and UniDiffuser is introduced after $2500$ iterations. In the second experiment, we generate samples from both PixArt-$\alpha$ and UniDiffuser, and a new prompt type is introduced after each $1000$ iterations. The results show that PAK-UCB-poly3 can well adapt to new prompt types and generators (see Figure LABEL:fig:varying_model).
@@
Setup 3. Synthetic experiments on other conditional generation tasks. We provide numerical results on image-captioning (image-to-text) and text-to-video (T2V) task under synthetic setups. The details can be found in Appendix D.
@@ 7 Conclusion
In this work, we investigated prompt-based selection of generative models using a contextual bandit algorithm, which can identify the best available generative model for a given text prompt. We propose two kernel-based algorithms, including PAK-UCB and RFF-UCB, to perform this selection task. Our numerical results on text-to-image, text-to-video, and image-captioning tasks demonstrate the effectiveness of the proposed framework in scenarios where the available generative models have varying performance rankings depending on the type of prompt. An interesting direction for future research is to extend the application of our algorithms to text-to-text language models, where different models may respond better to questions on different topics. Furthermore, considering evaluation criteria beyond relevance, such as diversity and novelty scores, could lead to extensions of our proposed framework.
@@
The technical challenge in analyzing PAK-UCB is that predictions in later iterations make use of previous outcomes. Hence, the rewards $\{s_{i}\}_{i\in\Psi_{g}}$ are not independent if the index set $\Phi_{g}$ is updated each time when model $g$ is chosen (line 6 of Algorithm 2). To address this problem, we leverage a standard approach used in prior works (Auer,, 2003; Chu et al.,, 2011; Valko et al.,, 2013) and present a variant of PAK-UCB in Algorithm 4, which is called Sup-PAK-UCB. We prove it attains a regret of order $\widetilde{O}(\sqrt{T})$.
A.1 The Sup-PAK-UCB Algorithm
@@
8:         Set upper confidence bound $\widehat{s}_{g}^{m}(y)\leftarrow\widehat{\mu}_{g}^{m}+(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}^{m}$ for all $g\in\widehat{{\mathcal{G}}}^{m}$.
9:         if $(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}^{m}\leq 1/\sqrt{T}$ for all $g\in\widehat{{\mathcal{G}}}^{m}$ then
@@
11:         else if $(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}^{m}\leq 2^{-m}$ for all $g\in\widehat{{\mathcal{G}}}^{m}$ then
12:              $\widehat{{\mathcal{G}}}^{m+1}\leftarrow\{g\in\widehat{{\mathcal{G}}}^{m}:\widehat{s}_{g}^{m}(y)\geq\max_{g\in\widehat{{\mathcal{G}}}^{m}}\widehat{s}_{g}^{m}(y)-2^{1-m}\}$.
@@
15:              Pick $g_{t}\in\widehat{{\mathcal{G}}}^{m}$ such that $(2\eta+\sqrt{\alpha})\cdot\widehat{\sigma}_{g}^{m}>2^{-m}$.
@@
A.2 Analysis
In this section, we prove the following regret bound of Sup-PAK-UCB.
@@ A.2 Analysis
Under Assumption 1, with probability at least $1-\delta$, the regret of Sup-PAK-UCB with $\eta=\sqrt{2\log(2(\log T)GT/\delta)}$ is bounded by
$\textup{Regret}(T)\leq\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)GT}\right)$ (6)
@@
Notations. To facilitate the analysis, we add an subscript $t$ to all the notations in Algorithm 4 to indicate they are quantities computed at the $t$-th iteration, i.e., $\widehat{\mu}_{g,t}^{m},\widehat{\sigma}_{g,t}^{m},\widehat{s}_{g,t}^{m},\widehat{{\mathcal{G}}}_{t}^{m}$, and $\Psi_{g,t}^{m}$. [^5] Note that line 14 in the algorithm is rewritten as “$\Psi_{g_{t},t+1}^{m}\leftarrow\Psi_{g_{t},t+1}^{m}\cup\{t\},\Psi_{g_{t},t+1}^{m^{\prime}}\leftarrow\Psi_{g_{t},t}^{m^{\prime}}$ for any $m^{\prime}\neq m$, and $\Psi_{g,t+1}^{m^{\prime}}\leftarrow\Psi_{g,t}^{m^{\prime}}$ for any $g\neq g_{t}$ and $m\in[M]$”. In addition, we set $\Psi_{g,t+1}^{m}\leftarrow\Psi_{g,t+1}^{m}$ for all $g\in{\mathcal{G}}$ and $m\in[M]$ in line 8.
We leverage the following two lemmas to prove Theorem 3. The first lemma shows that the construction of index sets $\{\Psi_{g,t}^{m}\}_{m=1}^{M}$ ensures the independence among the rewards $\{s_{i}\}_{i\in\Psi_{g,t}^{m}}$, which allows us to utilize Lemma 1 to bound the estimation error.
@@
The second lemma shows several properties of the estimated score $\widehat{s}_{g,t}^{m}$ and the set $\widehat{{\mathcal{G}}}^{m}_{t}$. The detailed proof can be found in Appendix C.2.
Lemma 4 (Valko et al., (2013), Lemma 7).
With probability at least $1-(MGT)\delta$, for any iteration $t\in[T]$ and stage $m\in[M]$, the following hold:
@@
Now, we are ready to finish the proof of Theorem 3.
Proof of Theorem 3.
1. Regret incurred in ${\mathcal{T}}_{0}$.
For any $t\in[T]$, let $m_{t}$ denote the stage that model $g_{t}$ is picked at the $t$-th iteration. We have that
$\sum_{t\in{\mathcal{T}}_{0}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))\leq$ $\sum_{t\in{\mathcal{T}}_{0}}(\widehat{s}_{g_{\star,t},t}^{m_{t}}(y)-s_{g_{t}}(y_{t}))$ (7)
$\leq$ $\sum_{t\in{\mathcal{T}}_{0}}(\widehat{s}_{g_{t},t}^{m_{t}}(y)-s_{g_{t}}(y_{t}))$
$=$ $\sum_{t\in{\mathcal{T}}_{0}}\left(\widehat{\mu}_{g_{t},t}^{m_{t}}+(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g_{t},t}^{m_{t}}-s_{g_{t}}(y_{t})\right)$
$\leq$ $2(2\eta+\sqrt{\alpha})\sum_{t\in{\mathcal{T}}_{0}}\widehat{\sigma}_{g_{t},t}^{m_{t}}$
$\leq$ $2(2\eta+\sqrt{\alpha})\sum_{t\in{\mathcal{T}}_{0}}T^{-\frac{1}{2}}\leq 2(2\eta+\sqrt{\alpha})\sqrt{T},$
where the second inequality holds by the definition of $g_{t}$ and the fact that $g_{\star,t}\in\widehat{{\mathcal{G}}}^{m_{t}}_{t}$, and the fifth inequality holds by line 7 of Algorithm 4.
2. Regret incurred in ${\mathcal{T}}_{1}$.
$\leq$ $\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}2^{3-m}\cdot|\Psi_{g,T+1}^{m}|,$
$(2\eta+\sqrt{\alpha})\sum_{t\in\Psi_{g,T+1}^{m}}\widehat{\sigma}_{g,t}^{m}>2^{-m}\cdot|\Psi_{g,T+1}^{m}|$
$\sum_{t\in\Psi_{g,T+1}^{m}}\widehat{\sigma}_{g,t}^{m}\leq\widetilde{O}\left(\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)|\Psi_{g,T+1}^{m}|}\right),$ (9)
$\sum_{t\in{\mathcal{T}}_{1}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))\leq$ $\widetilde{O}\left((1+\sqrt{\alpha})\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)|\Psi_{g,T+1}^{m}|}\right)$ (10)
$\leq$ $\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{GM}\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}|\Psi_{g,T+1}^{m}|}\right)$
$\leq$ $\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)GT}\right),$
where the second inequality holds by Cauchy-Schwarz inequality.
3. Putting everything together.
Combining Inequalities (7) and (10) leads to
$\textup{Regret}(T)=\left(\sum_{t\in{\mathcal{T}}_{0}}+\sum_{t\in{\mathcal{T}}_{1}}\right)(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))\leq\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{d_{\textup{eff}}\left(10+\frac{15}{\alpha}\right)GT}\right),$

### v3

@@ 6.1 Summary of the Numerical Results
The results of our numerical experiments indicate the improvement of the proposed PAK-UCB algorithm over the one-arm oracle baseline that is aware of the single generative model with the maximized averaged scores over the prompt distribution. This result suggests that the online learning algorithm could outperform a selector with the side-knowledge of the single best-performing model. We note that this improvement follows the prompt-based selection design of our proposed PAK-UCB algorithm. Moreover, our numerical results indicate that PAK-UCB can significantly improve upon the online learning baselines with shared weights, including LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) methods, indicating the effectiveness of the arm-specific design of PAK-UCB in the prompt-based model selection setting. Finally, in our experiments, the proposed RFF-UCB-RBF variant could reduce the computational costs of the general PAK-UCB algorithm.
Figure 4: Prompt-based selection of LLMs on code generation task (Setup 2): The Gemini-2.5-Flash-preview model attains higher pass rate on the code completion task while underperforms on the code translation task. Our proposed RFF-UCB-RBF algorithm queries solutions from the better model conditioned to the task category.
6.2 Experiment Settings
@@
Setup 1. Prompt-based selection among standard text-to-image models. The first set of experiments are on the setup illustrated in Figure 1(a), where we generate images from Stable Diffusion v1.5 and PixArt-$\alpha$. The results show that PAK-UCB-poly3 outperforms the baseline algorithm and attains a high optimal-pick-ratio, which shows that it can identify the optimal model conditioned to the prompt (see Figure LABEL:fig:t2itab1). Additionally, we provide numerical results on various T2I generative models, including UniDiffuser and DeepFloyd (see Figure LABEL:fig:t2itab2). Prompts are uniformly randomly selected from the MS-COCO dataset under two categories: ’dog’/’car’ (Figure LABEL:fig:t2itab1), ’train’/’baseball-bat’ (Figure LABEL:fig3a), ’elephant’/’fire-hydrant’ (Figure LABEL:fig3b), and ’carrot’/’bowl’ (Figure LABEL:fig3c). The input of the PAK-UCB, Naive-KRR, and RFF-UCB-RBF methods is the embedded prompt that is output by the pretrained CLIP-ViT-B-32-laion2B-e16 model from the open_clip repository [^7] https://github.com/mlfoundations/open_clip/tree/main. For LinUCB and KernelUCB baselines, we also concatenate the one-hot encoded vector of the model index to the CLIP-embedded prompt. Full results can be found in Figure LABEL:fig:496exp in the Appendix.
Setup 2. Prompt-based selection of LLMs. In these experiments, we focus on prompt-based selection of large language models (LLMs) on various tasks, including Sudoku-solving and code generation (code completion and translation). In the first setup, we select LLMs to solve code generation tasks. The model is given a Python code completion problem (sampled from the first $164$ tasks in BigCodeBench (Zhuo et al.,, 2024)) or C++-to-Java code translation problem (sampled from the HumanEval-X benchmark (Zheng et al.,, 2023)). Sample prompts can be found in Figure 13 and 14 in the Appendix. Particularly, the Gemini-2.5-Flash-preview model attains a higher pass@1 rate on code completion ($55.91$% versus $13.84$%) while underperforms for C++-to-Java translation ($24.39$% versus $40.91$%) compared to the Qwen-Plus model. The rewards are binary, indicating whether the generated code can pass all the test cases. The input of the PAK-UCB, Naive-KRR, and RFF-UCB-RBF methods is the $768$-dimensional embedded prompt that is output by the RoBERTa model (for code completion) or the CodeBERT model [^8] https://huggingface.co/microsoft/codebert-base (for code translation). For LinUCB and KernelUCB baselines, we also concatenate the one-hot encoded vector of the model index to the embedded prompt. The results (Figures 4(a) and LABEL:fig:code-gen-full) show that the proposed RFF-UCB-RBF method queries from the better model conditioned to the task category (Figure LABEL:fig:llm-selection).
In the second setup, we utilize two LLMs, including o3-mini and Deepseek-Chat, to solve $9\times 9$ Sudoku problems, which contains $5$-$30$ blank entries. Sample prompt can be found in Figure 11 in the Appendix. The o3-mini model attains a higher success rate on harder Sudoku problems with more than $15$ blank entries ($70.82$% versus $50.42$%), while slightly underperforms on easier Sudoku problems ($86.85$% versus $93.6$%). The rewards are binary, indicating the solution is correct or wrong. The input of the PAK-UCB, Naive-KRR, and RFF-UCB-RBF methods is the $768$-dimensional embedded prompt that is output by the RoBERTa model [^9] https://huggingface.co/docs/transformers/en/model_doc/roberta. For LinUCB and KernelUCB baselines, we also concatenate the one-hot encoded vector of the model index to the RoBERTa-embedded prompt. The results are summarized in Figures LABEL:fig:llm-selection(b) and LABEL:fig:sudoku-full, which show that the proposed RFF-UCB-RBF algorithm outperforms all the baseline methods.
@@
Setup 3. Adaptation to new prompt types and models. We consider scenarios where new text-to-image models or prompt types are introduced after the initial deployment. At the beginning of the first experiment, Stable Diffusion v1.5 and PixArt-$\alpha$ are available, and UniDiffuser is introduced after $2500$ iterations. Prompts are uniformly randomly selected from the MS-COCO dataset under categories ’train’ and ’baseball-bat’. In the second experiment, we generate samples from both PixArt-$\alpha$ and UniDiffuser. In the first 1k iterations, the prompts are uniformly randomly selected from a pool that initially includes categories ’person’ and ’bicycle’ in the MS-COCO dataset. Then, categories ’airplane’, ’bus’, ’train’, and ’truck’ are added to the pool after each 1k iterations. The results show that PAK-UCB-poly3 can well adapt to new prompt types and generators (see Figure LABEL:fig:varying_model). The input context follows Setup 1.
@@
Setup 4. Synthetic experiments on other conditional generation tasks. We provide numerical results on image-captioning (image-to-text) and text-to-video (T2V) task under synthetic setups in Appendix C.3.
@@ 7 Conclusion
In this work, we investigated prompt-based selection of generative models using a contextual bandit algorithm, which can identify the best available generative model for a given text prompt. We propose two kernel-based algorithms, including PAK-UCB and RFF-UCB, to perform this selection task. Our numerical results on LLMs (text-to-text), text-to-image, text-to-video, and image-captioning tasks demonstrate the effectiveness of the proposed framework in scenarios where the available generative models have varying performance rankings depending on the type of prompt. A potential future direction is to extend the online selection to capture the diversity and novelty of generated data. Also, developing cost-aware online selection methods, similar to the cost-aware CB algorithm in (Hu et al., 2025b, ), will be a relevant direction for future studies.
@@
The technical challenge in analyzing PAK-UCB is that predictions in later iterations make use of previous outcomes. Hence, the rewards $\{s_{i}\}_{i\in\Psi_{g}}$ are not independent if the index set $\Phi_{g}$ is updated each time when model $g$ is chosen (Line 6 of Algorithm 2). To address this problem, we leverage a standard approach used in prior works (Auer,, 2003; Chu et al.,, 2011; Valko et al.,, 2013) and present a variant of PAK-UCB in Algorithm 3, which is called Sup-PAK-UCB.
@@
8:         Set $\widehat{s}_{g}^{m}(y_{t})\leftarrow\widehat{\mu}_{g}^{m}+\eta\widehat{\sigma}_{g}^{m}$ and $\widetilde{s}_{g}^{m}(y_{t})\leftarrow\widehat{\mu}_{g}^{m}+(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g}^{m}$ for all $g\in\widehat{{\mathcal{G}}}^{m}$.
9:         if $(\eta+\sqrt{\alpha})\widehat{\sigma}_{g}^{m}\leq 1/\sqrt{T}$ for all $g\in\widehat{{\mathcal{G}}}^{m}$ then
@@
11:         else if $(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g}^{m}\leq 2^{1-m}$ for all $g\in\widehat{{\mathcal{G}}}^{m}$ then
12:              $\widehat{{\mathcal{G}}}^{m+1}\leftarrow\left\{g\in\widehat{{\mathcal{G}}}^{m}:\widetilde{s}_{g}^{m}(y_{t})\geq\max_{g\in\widehat{{\mathcal{G}}}^{m}}\widetilde{s}_{g}^{m}(y_{t})-2^{2-m}\right\}$.
@@
15:              Pick $g_{t}\in\widehat{{\mathcal{G}}}^{m}$ such that $(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g}^{m}>2^{1-m}$.
@@
A.1 Regret Analysis of Sup-PAK-UCB
@@
Under Assumption 1, with probability at least $1-\delta$, the regret of Sup-PAK-UCB with $\eta=\sqrt{2\log(2GT/\delta)}$ is bounded by
$\textup{Regret}(T)\leq\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{d_{\textup{eff}}\left(1+\frac{1}{\alpha}\right)GT}\right),$ (4)
@@
Notations. To facilitate the analysis, we add an subscript $t$ to all the notations in Algorithm 3 to indicate they are quantities computed at the $t$-th iteration, i.e., $\widehat{\mu}_{g,t}^{m}$ and $\widehat{\sigma}_{g,t}^{m}$ (first appeared Line 5),$\widehat{s}_{g,t}^{m}$ and $\widetilde{s}_{g,t}^{m}$ (first appeared Line 6), $\widehat{{\mathcal{G}}}_{t}^{m}$, and $\Psi_{g,t}^{m}$. In addition, we define $g_{\star,t}:=\operatorname*{arg\,max}_{g\in{\mathcal{G}}}s_{g}(y_{t})$ as the optimal model for prompt $y_{t}$ and $\widehat{g}_{\star,t}^{m}:=\operatorname*{arg\,max}_{g\in\widehat{{\mathcal{G}}}^{m}_{t}}\widehat{s}_{g,t}^{m}$ is optimistic model at stage $m$ [^10] Note that Line 14 of Algorithm 3 is rewritten as “$\Psi_{g_{t},t+1}^{m}\leftarrow\Psi_{g_{t},t+1}^{m}\cup\{t\},\Psi_{g_{t},t+1}^{m^{\prime}}\leftarrow\Psi_{g_{t},t}^{m^{\prime}}$ for any $m^{\prime}\neq m$, and $\Psi_{g,t+1}^{m^{\prime}}\leftarrow\Psi_{g,t}^{m^{\prime}}$ for any $g\neq g_{t}$ and $m\in[M]$”. In addition, we set $\Psi_{g,t+1}^{m}\leftarrow\Psi_{g,t+1}^{m}$ for all $g\in{\mathcal{G}}$ and $m\in[M]$ in Line 8.
Proof of Theorem 3.
1. Regret incurred in ${\mathcal{T}}_{0}$. For any $t\in[T]$, let $m_{t}$ denote the stage that model $g_{t}$ is picked at the $t$-th iteration. We have that
$\sum_{t\in{\mathcal{T}}_{0}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))=$ $\sum_{t\in{\mathcal{T}}_{0}}(s_{\star}(y_{t})-\widehat{s}_{g_{\star,t},t}^{m_{t}}(y)+\underbrace{\widehat{s}_{g_{\star,t},t}^{m_{t}}(y)-\widehat{s}_{g_{t}}^{m_{t}}(y_{t})}_{\text{$\leq 0$ by Line 8}}+\widehat{s}_{g_{t}}^{m_{t}}(y_{t})-s_{g_{t}}(y_{t}))$ (5)
$\leq$ $(\eta+\sqrt{\alpha})\sum_{t\in{\mathcal{T}}_{0}}\widehat{\sigma}_{g_{\star,t},t}^{m_{t}}+(3\eta+\sqrt{\alpha})\sum_{t\in{\mathcal{T}}_{0}}\widehat{\sigma}_{g_{t},t}^{m_{t}}$
$\leq$ $O(\sqrt{T}),$
where the first inequality holds by the fact that $g_{\star,t}\in\widehat{{\mathcal{G}}}^{m_{t}}_{t}$.
2. Regret incurred in ${\mathcal{T}}_{1}$. Note that
$\leq$ $\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}2^{3-m}\cdot|\Psi_{g,T+1}^{m}|,$
$\sum_{t\in\Psi_{g,T+1}^{m}}(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g,t}^{m}>2^{1-m}\cdot|\Psi_{g,T+1}^{m}|$
$\sum_{t\in\Psi_{g,T+1}^{m}}\widehat{\sigma}_{g,t}^{m}\leq\widetilde{O}\left(\sqrt{d_{\textup{eff}}\left(1+\frac{1}{\alpha}\right)|\Psi_{g,T+1}^{m}|}\right),$ (7)
$\sum_{t\in{\mathcal{T}}_{1}}(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))\leq$ $\widetilde{O}\left(\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}\sqrt{d_{\textup{eff}}\left(1+\frac{1}{\alpha}\right)|\Psi_{g,T+1}^{m}|}\right)$ (8)
$\leq$ $\widetilde{O}\left(\sqrt{GM}\sqrt{d_{\textup{eff}}\left(1+\frac{1}{\alpha}\right)\sum_{g\in{\mathcal{G}}}\sum_{m\in[M]}|\Psi_{g,T+1}^{m}|}\right)$
$\leq$ $\widetilde{O}\left(\sqrt{d_{\textup{eff}}\left(1+\frac{1}{\alpha}\right)GT}\right),$
where the second inequality holds by Cauchy-Schwarz inequality.
3. Putting everything together. Combining Inequalities (5) and (8) leads to
$\textup{Regret}(T)=\left(\sum_{t\in{\mathcal{T}}_{0}}+\sum_{t\in{\mathcal{T}}_{1}}\right)(s_{\star}(y_{t})-s_{g_{t}}(y_{t}))\leq\widetilde{O}\left((1+\sqrt{\alpha})\sqrt{d_{\textup{eff}}\left(1+\frac{1}{\alpha}\right)GT}\right),$
which concludes the proof. ∎
A.2 Auxiliary Lemmas
@@
Lemma 3.
With probability at least $1-(MGT)\delta$, for any iteration $t\in[T]$ and stage $m\in[M]$ in Algorithm 3, the following statements hold:
@@
Proof.
The first statement holds by both Lemma 2 and Lemma 4, and a uniform bound over all $t\in[T],g\in{\mathcal{G}}$, and $m\in[M]$.
To show the second statement, first note that $g_{\star,t}\in\widehat{{\mathcal{G}}}^{1}_{t}$. Assume $g_{\star,t}\in\widehat{{\mathcal{G}}}^{m}_{t}$ for some $m\in[M-1]$. Let $\widetilde{s}_{g,t}^{m}:=\widehat{\mu}_{g,t}^{m}+(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g,t}^{m}$ (computed in Line 10 in the algorithm). Then, by the first statement, we obtain that
$\widetilde{s}_{g_{\star,t},t}^{m}-\max_{g\in\widehat{{\mathcal{G}}}^{m}_{t}}\widetilde{s}_{g,t}^{m}=$ $\underbrace{\widetilde{s}_{g_{\star,t},t}^{m}-s_{\star}(y_{t})}_{\text{ $\geq 0$ }}+s_{\star}(y_{t})-\max_{g\in\widehat{{\mathcal{G}}}^{m}_{t}}\left\{\widetilde{s}_{g,t}^{m}-s_{g}(y_{t})+s_{g}(y_{t})\right\}$
$\geq$ $-\max_{g\in\widehat{{\mathcal{G}}}^{m}_{t}}\left\{\widetilde{s}_{g,t}^{m}-s_{g}(y_{t})\right\}$
$\geq$ $2(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g,t}^{m}\geq 2\cdot(-2^{1-m})=-2^{2-m},$
where the last inequality holds by Line 9 of the algorithm.
To show the last statement, note that for any $g\in\widehat{{\mathcal{G}}}^{m}_{t}$, it holds that
$s_{\star}(y_{t})-s_{g}(y_{t})=$ $\underbrace{s_{\star}(y_{t})-\widetilde{s}_{g_{\star,t},t}^{m}}_{\text{$\leq 0$}}+\widetilde{s}_{g_{\star,t},t}^{m}-\left(s_{g}(y_{t})-\widetilde{s}_{g,t}^{m}+\widetilde{s}_{g,t}^{m}\right)$
$\leq$ $2(2\eta+\sqrt{\alpha})\widehat{\sigma}_{g}^{m,t}+|\widetilde{s}_{g_{\star,t},t}^{m}-\widetilde{s}_{g,t}^{m}|$
$\leq$ $2\cdot 2^{1-m}+2^{2-m}\leq 2^{3-m},$